OpenStudy (goformit100):

A particle undergoes two displacements. The first has a magnitude of 150 cm and makes an angle of 120° with the positive x axis. The resultant displacement has a magnitude of 140 cm and is directed at an angle of 35.0° to the positive x axis. Find the magnitude and direction of the second displacement.

OpenStudy (goformit100):

@bahrom7893

OpenStudy (bahrom7893):

god i hate vectors. @.Sam.

sam (.sam.):

Hmm alright I'll think of it since you're going for it too

OpenStudy (bahrom7893):

Ok, this is what I found: x component of the total final displacement = 140 cos 35 = 114.68 y component of the total final displacement = 140 sin 35 = 80.3 x component of the first displacement vector = -150 x cos 60 = -75 y component of the first displacement vector = 150 x sin 30 = 75 Thus, x component of the second displacement vector = 114.68 - 75 = 39.68 y component of the second displacement vector = 80.3 - (- 75) = 155.3 Magnitude of the 2nd vector is: sqrt(39.68^2+155.3^2) = 160.28

OpenStudy (bahrom7893):

trying to draw the picture now http://www.twiddla.com/1166652

sam (.sam.):

Erm \[Y:150\cos(30)+xsin(\theta)+140\sin(35)=0 \\ \\ X:150\sin(30)=xcos(\theta)+140\sin(35)\]